# A sobel-gradient MLP baseline for handwritten character recognition

### v1

@@ Abstract
keywords:
handwritten recognition , Sobel operator , edge detection , multilayer perceptron , MNIST , EMNIST
journal: Pattern Recognition Letters
[…]
[a]organization=Department of Science, Technology & Mathematics, Lincoln University, addressline=820 Chestnut Street, city=Jefferson City, state=MO, postcode=65101, country=USA
@@ 2 Related Work
Edge-aware CNNs
@@
All-MLP architectures
Token-mixing MLPs (e.g., MLP-Mixer [8]) show dense models can be competitive when inputs are suitably structured. We push this to an extreme: only two derivative channels.
@@ 3.1 Intuitive Explanation: Why Sobel Detects Edges
Why does this matter for handwritten glyphs? Because the strokes are only a pixel or two thick; almost all the information that distinguishes a C from a G lies in these sharp intensity changes. By feeding the raw Sobel maps to an MLP, we give the network exactly those edge patterns without any extra noise or texture.
@@ 4 Method
$\widehat{G}_{x}$ $=\frac{G_{x}-\min(G_{x})}{\max(G_{x})-\min(G_{x})+\varepsilon},$ (1)
$\widehat{G}_{y}$ $=\frac{G_{y}-\min(G_{y})}{\max(G_{y})-\min(G_{y})+\varepsilon},\qquad\varepsilon=10^{-8}.$ (2)
@@
Parameter count (exact)
@@ 5 Experimental Setup
Split protocol
We concatenate TFDS train and test and perform a stratified 80/20 split using sklearn.model_selection.train_test_split.
During training, Keras holds out 10% of the training portion as validation via $\texttt{validation\_split}=0.1$. Table 2 shows the resulting sizes.
@@
width Dataset Total Train (fit) Val (held-out) Test MNIST (0–9) 70,000 50,400 5,600 14,000 EMNIST Letters (A–Z) 145,600 104,832 11,648 29,120
Label handling
@@
Evaluation metrics and class balance
@@
Environment
OpenCV for Sobel (cv2.Sobel); TensorFlow/Keras 2.x; TFDS; NumPy; scikit-learn.
@@ 7 Discussion
First-order edges capture core glyph structure
@@ 8 Conclusion
https://github.com/MN-21/Sobel-Gradient-MLP.

### v2

@@ Abstract
Keywords: handwritten recognition; Sobel operator; edge detection; multilayer perceptron; MNIST; EMNIST
@@ 2 Related Work
Edge-aware CNNs.
@@
All-MLP architectures.
Token-mixing MLPs (e.g., MLP-Mixer [10]) show dense models can be competitive when inputs are suitably structured. We push this to an extreme: only two derivative channels. Relatedly, curvature-based descriptors have also been explored [4].
@@ 3.1 Intuitive Explanation: Why Sobel Detects Edges
Why does this matter for a handwritten glyphs? Because the strokes are only a pixel or two thick; almost all the information that distinguishes a C from a G lies in these sharp intensity changes. By feeding the raw Sobel maps to an MLP, we give the network exactly those edge patterns without any extra noise or texture.
@@ 4 Method
$\widehat{G}_{x}\;=\;\frac{G_{x}-\min(G_{x})}{\max(G_{x})-\min(G_{x})+\varepsilon},\qquad\widehat{G}_{y}\;=\;\frac{G_{y}-\min(G_{y})}{\max(G_{y})-\min(G_{y})+\varepsilon},\quad\varepsilon=10^{-8}.$
@@
Parameter count (exact).
@@ 5 Experimental Setup
Split protocol.
We concatenate TFDS train and test and perform a stratified 80/20 split using sklearn.model_selection.train_test_split. During training, Keras holds out 10% of the training portion as validation via $\texttt{validation\_split}=0.1$. Table 1 shows the resulting sizes.
@@
Dataset Total Train (fit) Val (held-out) Test
MNIST (0–9) 70,000 50,400 5,600 14,000
EMNIST Letters (A–Z) 145,600 104,832 11,648 29,120
Label handling.
@@
Evaluation metrics and class balance.
@@
Environment.
OpenCV for Sobel ($\texttt{cv2.Sobel}(\cdot,\cdot,1,0,\texttt{ksize}=3)$ and $\texttt{cv2.Sobel}(\cdot,\cdot,0,1,\texttt{ksize}=3)$); TensorFlow/Keras 2.x; TFDS; NumPy; scikit-learn.
@@ 7 Discussion
First-order edges capture core glyph structure.
@@ 8 Conclusion
https://github.com/AzamNouri/Sobel-Gradient-MLP.
For reproducibility, an archived release is preserved on Zenodo [5].